1.3 The monthly salaries of the employees at Lyman Limited follow a normal distribution with an average salary of R 14 000 and a standard deviation of R 500.

a) What percentage of employees get paid between R 13 000 and R 15 500?


1
Expert's answer
2021-09-28T11:44:13-0400

"\mu = 14000 \\

\sigma=500 \\

P(13000<X<15500) = P(X<15500) -P(X<13000) \\

=P(Z< \frac{15500-14000}{500}) -P(Z< \frac{13000-14000}{500}) \\

= P(Z< 3) -P(Z< -2) \\

= 0.9986 -0.0227 \\

= 0.9759 \\

= 97.59 \; \%"